the value of lim a^n+b^n/a^n-b^n,n->infinity,(where a>b>1)?

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

let f(n)=(a^n +b^n)/(a^n-b^n)=((a/b)n+1)/((a/b)n-1), it is easy to ee that as n→∞, (a/b)^n →∞,and f(n)→((a/b)n/(a/b)n)=1